The conjugate of 5 + 6i is
Find the product (2 + i)(3 + 4i). Give your answer in the usual a + bi form
Write the complex number i 15 in the usual a + bi form
Find the reciprocal of the complex number 3 + 2i . Give your answer in the usual a + bi form.
Write the complex number in the usual a + bi form.
What are the coordinates of the complex number -i when we plot it in the complex plane?
Find the magnitude of the complex number 3 - 7i. (In other words, find | 3 - 7i | )

Answer by stanbon(48545)

The conjugate of 5 + 6i is 5-6i
---------------
Find the product (2 + i)(3 + 4i). Give your answer in the usual a + bi form
Use FOIL:
F:2*3=6
O:2*4i=8i
I:3*i=3i
L:i*4i=-4
Combine to get:2+11i
-----------------------
Write the complex number i 15 in the usual a + bi form
0+15i
----------------
Find the reciprocal of the complex number 3 + 2i . Give your answer in the usual a + bi form.
1/(3+2i)=[3-2i)}/[3+2i)(3-2i)]=[ 3-2i]/[13]=(3/13)-(2/13)i
------------------
What are the coordinates of the complex number -i when we plot it in the complex plane?
(0,-1)
---------------
Find the magnitude of the complex number 3 - 7i. (In other words, find | 3 - 7i | )
|3-7i|=sqrt(9+49)=sqrt58
Cheers,
Stan H.